Can't we do the Particular solution by assuming sin(z) as imaginery part of e^iz,plz can u show how to do by that method


1
Expert's answer
2021-05-24T15:27:10-0400

Yes , we can find particular solution by assuming sin(z) as imaginary part of "e^{iZ}" .


let take a function Z with a variable t , so we write this function Z = cos t + i sin t


dz = (-sint + i cost )dt,


Taking i common from whole function , we get ,


dz = i(cost + i sint)dt = izdt


now seperating the variables , we get

-

"\dfrac{dz}{z}=idt" ,


now integerating on both side of the function , we get -


"\int\dfrac{dz}{z}=\int idt"


ln z = it ,


now removing the function ln ,


Z="e^{it}" +c,


so this is a particular solution when we assume sin z as imaginary part of "e^{iz}."